OpenStudy (anonymous):

What is an equation for the line that passes through the coordinates (-1,2) and (7,6) ?

hero (hero):

Using y = mx + b, insert the points as follows: 2 = -1m + b 6 = 7m + b Now isolate b: b = 2 + m b = 6 - 7m Next set b = b, then isolate and solve for m: 2 + m = 6 - 7m Place like terms on the same side: 7m + m = 6 - 2 8m = 4 m = 4/8 m = 1/2 Now find the value of b using the value found for m: b = 2 + m b = 2 + 1/2 b = 4/2 + 1/2 b = 5/2 Therefore the equation of the line is: y = x/2 + 5/2 y = (x + 5)/2

OpenStudy (anonymous):

did @hero help you or do you still not understand

hero (hero):

What is there to not understand? Insert the points, solve for the variables, find their values, then insert the values into the equation of the line.